### v1

# GeoTransformer: Enhancing Urban Forecasting with Geospatial Attention Mechanisms

Abstract
Recent advancements have focused on encoding urban spatial information into high-dimensional spaces, with notable efforts dedicated to integrating sociodemographic data and satellite imagery. These efforts have established foundational models in this field. However, the effective utilization of these spatial representations for urban forecasting applications remains under-explored. To address this gap, we introduce GeoTransformer, a novel structure that synergizes the Transformer architecture with geospatial statistics prior. GeoTransformer employs an innovative geospatial attention mechanism to incorporate extensive urban information and spatial dependencies into a unified predictive model. Specifically, we compute geospatial weighted attention scores between the target region and surrounding regions and leverage the integrated urban information for predictions. Extensive experiments on GDP and ride-share demand prediction tasks demonstrate that GeoTransformer significantly outperforms existing baseline models, showcasing its potential to enhance urban forecasting tasks.
Introduction
In urban forecasting tasks involving predicting economic indicators and human mobility, classical methods usually rely on low-dimensional numeric data, which includes Point of Interest (POI) data(Li et al. 2022), survey data(Roskam and Kunst 2008; DHS et al. 2013), GPS records (Wang et al. 2023a; Moreira-Matias et al. 2013; Pappalardo et al. 2013), demographic census, geospatial features and so forth.
Recently, many researchers have explored using high-dimensional information to fully capture the complexities of urban dynamics. A main direction is utilizing urban imagery for feature extraction and prediction (Ayush et al. 2020; Li et al. 2022; Yeh et al. 2020; Abdelhalim and Zhao 2024). A more advanced direction is encoding the urban information to a high-dimensional space representation, either through spatial representation learning (Mai, Li, and Lao 2023; Wu et al. 2024) or fusing satellite images and sociodemographic data (Wang et al. 2024). While these studies have established frameworks for encoding urban information, there has been limited investigation into optimizing these high-dimensional data for urban forecasting.
Numerous established methods in natural language processing (NLP) and computer vision have effectively harnessed high-dimensional data. Particularly, Transformer architecture and attention mechanisms have proven highly effective in decoding high-dimensional representations of natural language (Vaswani et al. 2017a; Devlin et al. 2018; Radford et al. 2018) and images (Dosovitskiy et al. 2020a; Rombach et al. 2022). Urban representations, characterized by high dimensionality and spatial correlations, are well-suited for applying similar decoding methods. Some studies have explored the use of attention mechanisms to capture spatial dependencies, particularly through Graph Attention Networks (GAT) (Feng and Tassiulas 2022; Zheng et al. 2020). However, these approaches primarily depend on graph structures and spatial embeddings. In fact, incorporating geospatial statistical priors might be more effective in encoding spatial dependencies, potentially reducing the need for complex data structuring and simplifying the system design.
To address this issue, we propose GeoTransformer, a deep architecture akin to Transformer (Vaswani et al. 2017b) that combines geospatial characteristics prior to the attention mechanism. The novel geospatial attention mechanism calculates attention between different urban regions weighted by distance, incorporating information from surrounding areas. This approach implicitly includes spatial dependencies in the attention mechanism. Extensive experiments demonstrate that our model significantly outperforms several baseline methods in predicting GDP and ride-share demand.
Our main contributions can be summarized as follows:
  • We propose GeoTransformer, a transformer-akin deep architecture designed for decoding high-dimensional urban representations and enhancing prediction accuracy.
  • We introduce a geospatial attention mechanism that prioritizes attention between regions weighted by spatial proximity, assigning higher weights to closer regions.
  • Experiments demonstrate that our method significantly outperforms baseline models in tasks of GDP and ride-share demand prediction.
Related Work
High Dimensional Input for Urban Forecasting
Urban forecasting leverages multiple data sources and advanced computational techniques to address challenges faced by cities. Traditional methods usually adopt low-dimensional numeric data for prediction. In recent years, with the need rising for complex prediction tasks and the development of deep learning techniques, researchers have explored various ways to leverage high-dimensional urban information. One direction is directly using various urban imagery for feature extraction and prediction. Albert, Kaur, and Gonzalez employ Convolutional Neural Networks (CNNs) to detect land-use patterns from satellite images. Ayush et al.; Li et al.; Yeh et al. utilize urban imagery to predict socioeconomic status. Abdelhalim and Zhao leverage roadside imagery to predict transit travel time.
Recently, encoding urban information into a high-dimensional representation marks a new direction in enhancing flexibility and prediction accuracy for different urban forecasting tasks. Mai, Li, and Lao develop spatial representation learning (SRL), utilizing deep neural networks (DNNs) to encode and feature spatial data in the forms of points, polylines, polygons, graphs, etc. Wang et al. propose a theoretical framework deep hybrid model (DHM) fusing demographic numeric data and satellite images into urban latent representations. However, how to leverage these urban representations more effectively in prediction tasks remains unexplored.
Transformer and attention mechanism(Vaswani 2017) has been widely adopted for its efficiency in capturing relationships within high-dimensional representations. They are first proposed to solve natural language understanding and then visual models like ViT(Dosovitskiy et al. 2020b) and Stable Diffusion (Rombach et al. 2022) also utilize attention mechanisms to comprehend images and the spatial relationships between image patches. Given its strength in handling high-dimensional data and capturing spatial relationships, the attention mechanism is well-suited for adaptation in modeling urban high-dimensional representations.
Approaches to Modeling Spatial Dependency
Spatial dependency has traditionally been addressed using two primary approaches. First, feature engineering methods like k-nearest neighbors (KNN) are employed to convert surrounding spatial information into features. Second, spatial statistical methods such as spatial error regression, spatial lag models(Anselin 2013), and geographically weighted regression (GWR) (Fotheringham, Charlton, and Brunsdon 1998) have been widely used to account for spatial autocorrelation.
In recent years, Graph Neural Networks (GNNs) have become a widely adopted method for capturing spatial dependencies, especially in spatiotemporal prediction(Wang et al. 2023b; Zhuang et al. 2023; Cai et al. 2020). Urban built environment data such as road networks, are inherently suitable for representation as graph structures due to their topological properties. However, grid-based high-dimensional urban representations often lack clear graph-like structures, making graph construction particularly challenging. Moreover, the intrinsic complexity of such high-dimensional data can further hinder the effective training of models.
Recent advancements have seen the integration of graph structures with attention mechanisms (Zheng et al. 2020; Feng and Tassiulas 2022), particularly through Graph Attention Networks (GATs)(Zhang, Yu, and Liu 2019). While these models incorporate spatiotemporal attention mechanisms, their reliance on graph structures and spatial embeddings remains significant for capturing spatial dependencies. Integrating spatial statistical priors could simplify the model’s reliance on complex data structures and streamline system design. Tobler’s First Law of Geography (Tobler 1970), which posits that closer entities are more likely to be related, could provide a valuable framework for capturing spatial dependency. Despite its potential, the integration of spatial statistical priors with attention mechanisms remains unexplored.
Therefore, in this study, we aim to introduce an attention mechanism guided by spatial statistical priors as a new paradigm for effectively leveraging high-dimensional urban representation data.
Economic and Mobility Indicators Prediction
Gross Domestic Product (GDP) is a critical indicator in urban studies, offering vital insights into the economic health, developmental patterns, and sustainability of cities(Ribeiro et al. 2021; Mahtta et al. 2022; Brilhante and Klaas 2018). Leveraging Nighttime Light (NTL) data to estimate GDP as an economic indicator has become a widely adopted approach (Liu et al. 2021; Sun et al. 2020). There is substantial research indicating a strong association between GDP and urban spatial characteristics (Zhang, He, and Yuan 2020; Zhaoa and Chenb 2023), with numerous studies employing machine learning methods to model and predict regional GDP using structured urban information or urban imagery data(Kopczewska 2022; Chen et al. 2022b). For example, Chen et al. utilize deep learning methods to predict county-level GDP grids. While substantial research has established the relationship between GDP and urban spatial characteristics, the application of high-dimensional urban representations for GDP prediction remains underexplored. Given the critical role of GDP in validating urban prediction models, we selected GDP estimates derived from Nighttime Light (NTL) data as a benchmark for evaluating our approach.
In addition to economic indicators, mobility indicators such as ride-share demand serve as a critical indicator for urban forecasting, and it has significant implications for transportation planning, economic activity, and population behavior (Komanduri et al. 2018; Ma and Hanrahan 2020; Zhang and Zhang 2018). For instance, Shaheen et al. emphasize that ride-share services can influence traffic patterns, reduce private vehicle ownership, and improve accessibility in urban environments. Researchers have developed utilizing deep learning methods to predict ride-share demand from urban spatial information(Mühlematter et al. 2023; Zhao et al. 2022; Zhang and Zhao 2022). Zhao et al. couple graph neural networks (GNNs) with the spatial-temporal influence of the built environment. Zhang and Zhao propose a Clustering-aided Ensemble Model to improve ridesourcing demand forecasting. Similarly, few studies have explored using high-dimensional urban representations to predict ride-share demand. Given the significance of ride-share demand as a key indicator with strong spatial dependencies, it serves as an ideal benchmark for evaluating the efficacy of our proposed method.
Methodology
As illustrated in Figure 1, our framework adopts an encoder-decoder structure. The encoder leverages existing state-of-the-art methods to process numeric and imagery inputs, representing city regions in a latent space through a mixing operator. Our novel GeoTransformer decoder model then utilizes this latent representation to predict GDP and ride-share demand.
Figure 1: Our urban forecasting framework uses a mixing operator to fuse information into latent representations which are then passed to GeoTransformer for prediction.
Encode the city: Mixing Operator
To fuse satellite imagery and sociodemographic information into a high-dimensional latent representation, we adopt a mixing operator $\mathcal{M}$, which can be represented as:
$z_{n}=\mathcal{M}(x_{n},I_{n})$ (1)
in which $x_{n}$ and $I_{n}$ represent the sociodemographics and satellite imagery inputs, and $z_{n}$ represents the latent representation of an urban area.
Following the work Deep Hybrid Model (Wang et al. 2024), we adopt a supervised autoencoder (SAE). The SAE maps the imagery into latent variables through the encoder $E(I_{n})$, and reconstructs images through the decoder $D(z_{n})$ and sociodemographic through a supervision network $F(z_{n})$ (Figure 2 ). It is trained by optimizing the sum of autoencoder loss $\mathcal{L}_{AE}$ and sociodemographic supervision loss $\mathcal{L}_{sup}$:
$\mathcal{L}_{SAE}=(1-\lambda)\mathcal{L}_{AE}+\lambda\mathcal{L}_{sup}$ (2)
where $\mathcal{L}_{sup}$ denotes the absolute error of the sociodemographic variables $x$ $\mathcal{L}_{AE}$ represents the sum of L1 image reconstruction loss $\mathcal{L}_{rec}$, the KL divergence loss $\mathcal{L}_{kl}$, and the learned perceptual image patch similarity loss $\mathcal{L}_{lpipis}$:
$\mathcal{L}_{sup}=\alpha_{0}|x-\hat{x}|$ (3)
$\mathcal{L}_{AE}=\alpha_{1}\mathcal{L}_{rec}+\alpha_{2}\mathcal{L}_{kl}+\alpha_{3}\mathcal{L}_{lpipis}$ (4)
in which all the $\alpha$ terms are the adaptive hyperparameters used for scaling each loss.
For each autoencoder loss, $\mathcal{L}_{rec}$ measures the L1 distance between the input images $I_{n}$ and reconstructed images $\hat{I_{n}}$ in the pixel space, $\mathcal{L}_{kl}$ measures the distance between the estimated latent distribution and the standard normal prior distribution, and $\mathcal{L}_{lpips}$ is calculated by extracting feature stacks from $L$ layers of a pre-trained neural network, unit-normalizing them in the channel dimension ($\hat{y}^{l}_{I_{n,hw}},\hat{y}^{l}_{\hat{I}_{n,hw}}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$), and scaling by vector $w_{l}$ before computing the L2 distance.
The mixing operator allows for the fusion of sociodemographics and satellite imagery information in a high-dimensional latent space. Each region of a city is then encoded into a latent representation using the encoder of the trained SAE.
Figure 2: The mixing operator is trained by extracting sociodemographic information and reconstructing the input satellite images from latent space.
GeoTransformer: Geospatial-weighted Attention Mechanisms
We propose a novel approach GeoTransformer to decode the urban latent representations for downstream tasks including GDP and ride-share demand prediction.
As shown in Figure 3, our method adopts a structure akin to the Transformer (Vaswani 2017) decoder. Since the downstream tasks do not involve sequence prediction, the self-attention module is optional in this study. We introduce a geospatial attention mechanism to capture the spatial dependencies and provide location information between city regions. The outputs of the decoder layers are then passed through a fully connected linear layer. The whole system $\mathcal{G}$ can be represented as:
$y_{n}=\mathcal{G}(z_{n},\text{kNN}(z_{n},Z))$ (5)
in which the input includes the latent representation of a region $z_{n}$ and its k-nearest neighbors’ latent representations $\text{kNN}(z_{n},Z)$ from all regions $Z$, and $y_{n}$ is the predicted output.
The geospatial attention mechanism is the core component of our method. It computes attention scores between city regions and applies geospatial weighting to these scores. The attention calculation is represented as:
$\text{GeoAtt}(Q,K,V)=W\odot\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (6)
where the query matrix $Q$ contains the latent representation of the query region $z_{n}$ for which the target variable is being predicted, while the value matrix $V$ consists of $z_{n}$ itself and the latent representations of its neighbors as value regions. $d_{k}$ denotes the dimension of the key vectors. Intuitively, as the similarity between two city regions is not directly related to target variables, we set the key matrix $K$ as trainable weights rather than being identical to $V$.
$W$ is the spatial weighting factor, computed based on the spatial distance between query regions in $Q$ and value regions in $V$. Our model provides linear and non-linear spatial weighting options based on the assumption of Tobler’s First Law of Geography (Tobler 1970), suggesting that regions closer to the query region are assigned higher weights.
Linear Normalization
scales the distances within 0 to 1, inversely transforming shorter distances into higher weights:
$W_{j}=1-\frac{d_{j}}{\text{max}(d)}$ (7)
where $d_{j}$ is the distance to the $j^{th}$ value region, and $\text{max}(d)$ are the maximum distances observed.
Inverse Distance Weighting (IDW) (Shepard 1968)
calculates weights inversely proportional to the distance:
$W_{j}=\frac{1}{d^{p}_{j}+\epsilon}$ (8)
where $d_{j}$ is the distance to the $j^{th}$ value region, the power value $p$ is set to 2 in standard applications, and $\epsilon$ is a small constant to prevent division by zero. The weights are then subsequently normalized to ensure they sum to unity.
Gaussian Weighting
applies a Gaussian function to the distances, providing a gradual decrease in weight as distance increases:
$W_{j}=e^{-(\frac{d^{2}_{j}}{2\sigma^{2}})}$ (9)
where $d_{j}$ is the distance to the $j^{th}$ value region, and $\sigma$ determines the spread of the weights. The weights are then subsequently normalized to ensure they sum to unity.
Adhering to the established multi-head attention mechanism (Vaswani 2017), our model employs multiple attention heads to capture various spatial relationships concurrently. The calculation can be represented as:
$\text{MultiHead}(Q,K,V)=\text{Concat}(head_{1},...,head_{h})W^{O}$ (10)
$\textbf{where }\text{head}_{i}=\text{GeoAtt}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i})$ (11)
where $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$ and $W^{O}$ are the projection matrices. The amount of heads and layers of the decoder module is also adaptive.
The GeoTransformer architecture is designed to extract geospatial information in high-dimensional latent space, and it is also adaptable to various geospatial prediction tasks due to the flexibility of the output layers. The design of geospatial attention allows the model to capture and integrate spatial dependencies, utilizing local context from both the target region and its surroundings for more accurate predictions. Leveraging high-dimensional urban information and deep architectures, our GeoTransformer represents a novel approach to geospatial analysis and urban planning applications.
Figure 3: The GeoTransformer structure includes a geospatial attention module that calculates the geospatial weighted attention scores between the query region and its neighboring value regions.
Panel 1 Baselines with only numeric or imagery input
Model 1: Linear Regression sociodemographic numeric input (12 dim)
Model 2: ViT satellite imagery input (4*512*512 dim), heads = 12, layers = 12
Panel 2 Baselines with urban latent representations input
Model 3: Polynomial Regression latent input (reduced dim from 4096 to 128 using PCA), poly degree = 2, $l_{2}$ regularization with $\lambda$ = 0.5
Model 4: GAT latent input (4096 dim), layers = 4
Panel 3 GeoTransformer with 3 geospatial weighting methods
Model 5: Linear Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=49 / 81
Model 6: IDW Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=49 / 81
Model 7: Gaussian Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=81 / 169
Table 1: Training settings of baselines and GeoTransformer models.
GDP and Ride-share Demand Modeling
For the GDP modeling, we use the Global 1km×1km gridded revised data (Chen et al. 2022a) which is based on night-time light data from DMSP/OLS and NPP/VIIRS sensors. The data is calibrated across different years and sensors to produce a consistent high-resolution dataset. A regression model is constructed by linking night-time light intensity, population density, and other auxiliary variables with actual GDP. the gridded real GDP can be calculated using:
$RGY_{ij,t}^{*}=\begin{cases}\frac{RGY_{ij,t+1}^{*}}{1+gy_{ij,t}},&\text{if }DN_{ij,t}\neq 0\\ 0,&\text{if }DN_{ij,t}=0\end{cases}$ (12)
where $RGY_{ij,t}^{*}$ denotes the $j^{th}$ grid in the $i^{th}$ country’s real GDP in the period of $t$ based on the revised real growth rate. The calculations are based on the hypothesis that there is no GDP when the $DN$ value is zero. The original GDP data, gridded at 1km×1km resolution, is resampled using bilinear interpolation.
For the ride-share demand modeling, we utilize taxi trip data from Chicago that is constant with the time frame of the remote sensing imagery. Focusing on the pickup locations to represent ride-share demand, we aggregate the demand for each census tract by calculating the total number of pickups within each tract. Subsequently, we apply area-weighted averaging to estimate the ride-share demand corresponding to each imagery. The area-weighted average ride-share demand $\overline{D}$ for each imagery is calculated as follows:
$\overline{D}=\frac{\sum_{i=1}^{n}D_{i}\cdot A_{i}}{\sum_{i=1}^{n}A_{i}}$ (13)
where $D_{i}$ is the ride-share demand (number of pickups) in the $i$-th census tract, $A_{i}$ is the area of the $i$-th census tract, and $n$ is the total number of census tracts. This approach allows us to align the spatial resolution of the taxi trip data with the remote sensing imagery, facilitating accurate and effective modeling of ride-share demand.
Model GDP Ride-share Demand
MSE $\downarrow$ MAE $\downarrow$ $R^{2}$ MSE×$\times 10^{7}$downarrow MAE×$\times 10^{3}$downarrow $R^{2}$
Panel 1 Baselines with only numeric or imagery input
1 Linear Regression 3.22/3.13 1.44/1.43 0.159/0.151 8.45/6.11 8.36/5.99 0.668/0.720
2 ViT 0.616/0.828 0.595/0.703 0.822/0.759 15.6/22.5 4.10/4.16 0.526/0.315
Panel 2 Baselines with urban latent representations input
3 Polynomial Regression 2.19/3.25 1.17/1.46 0.429/0.118 12.8/19.5 5.05/6.00 0.494/0.335
4 GAT 0.469/0.726 0.476/0.602 0.866/0.805 7.21/9.51 2.01/2.35 0.750/0.575
Panel 3 GeoTransformer with 3 geospatial weighting methods
5 Linear Weighting 0.081/0.216 0.210/0.346 0.979/0.944 1.62/7.26 0.98/2.53 0.934/0.779
6 IDW Weighting 0.570/0.686 0.537/0.585 0.849/0.822 0.172/3.02 0.491/1.16 0.993/0.908
7 Gaussian Weighting 0.448/0.567 0.486/0.533 0.882/0.852 0.263/9.78 0.469/3.97 0.989/0.703
Table 2: Predictive performance of baselines and GeoTransformer models. Each entry is represented as training/testing performance.
Experiment
In this section, we provide our method’s results in GDP and ride-share demand prediction and compare them with baseline methods. Ablation experiments are conducted to demonstrate the effectiveness of the geospatial attention component and model architecture design.
Experiment Setup
Data Preparation.
For the satellite imagery data, we utilized the National Agriculture Imagery Program (NAIP) four-band remote sensing imagery for the Greater Chicago Area, acquired in September 2019. The imagery has a high spatial resolution of 0.6 meters. Using Google Earth Engine, we download the dataset and subsequently divide it into tiles of size 512x512 to facilitate our analysis.
For spatial correspondence between structured and unstructured information, we created a fishnet grid with a cell size of 500m×500m, aligned it with the same geospatial location as the maptile data, and resampled the 1km×1km gridded GDP data to match the fishnet.
For the ride-share demand, we use the Taxi Trips dataset obtained from the City of Chicago. The dataset includes fundamental trip information such as unique ID, pickup and drop-off timestamps, latitude and longitude coordinates, trip miles, fare, and other relevant details. To align the spatial resolution of the taxi trip data with the satellite imagery tiles, we performed a spatial join. This process allows us to calculate the ride-share demand corresponding to each tile, ensuring accurate and effective integration of the two data sources.
In addition to the imagery and ride-share data, we incorporate sociodemographic data from the American Community Survey (ACS) 2015-2019. This dataset provides detailed information at the census tract level, including variables such as median rent, racial composition, median household income, the Gini index, and so on.
Mixing Operator Training.
We utilize the VAE component from Stable Diffusion (Rombach et al. 2022) as the initial weights for the mixing operator. To accommodate a 4096-dimensional latent space, linear layers are introduced at both the encoder’s output and the decoder’s input. The decoder reconstructs images using mean absolute error (MAE) as the reconstruction loss. An annealing strategy is adopted for the KL divergence loss, gradually increasing its weight ratio. Additionally, the latent space restores sociodemographic information through a feed-forward module, employing mean squared error (MSE) as the loss metric. The training process spans 100 epochs in total. The latent representation for each region is inferred using the encoder part of the mixing operator.
Baselines.
We train various baseline models for comparison. As detailed in Table 1, models in Panel 1 utilize either direct numeric or imagery input. Specifically, Model 1 employs a linear regression configured with 12-dimensional sociodemographic data as input. Model 2 utilizes a pre-trained Visual Transformer (ViT) (Dosovitskiy et al. 2020a), which is subsequently fine-tuned on the prediction tasks with the last 4 encoder layers and the output layer unfrozen. Panel 2 presents models that utilize latent representations as inputs to perform predictions. Following DHM(Wang et al. 2024), model 3 employs a polynomial regression. Model 4 adopts a Graph Attention Network (GAT) (Zhang, Yu, and Liu 2019) model with each city region as a node connected to its neighbors, forming the graph structure. All hyperparameters for training the baseline models are optimized through a grid search to ensure peak performance on the test set.
GeoTransformer Training.
As shown in Table 1 panel 3, we train the GeoTransformer (model 5-7) using three spatial weighting methods. The number of decoder layers is set to 4, and the number of attention heads is set to 16. For each model variant and task, the L2 regularization decay weight was optimized, with a range of 0.0001 to 0.1, while the number of neighboring regions is set to 49, 81, or 169. All parameters are grid-searched to ensure the best performance. For both GDP and ride-share demand predictions, mean squared error (MSE) is employed as the loss function. All models are trained on an RTX 4090 GPU running Ubuntu 22.04.
Evaluation Metrics.
We employ commonly used statistical metrics that evaluate the accuracy of the predictions, including Mean Squared Error (MSE), Mean Absolute Error (MAE), and R-squared ($R^{2}$). MSE evaluates the average squared differences between predicted and actual values, MAE measures the average absolute differences, and ($R^{2}$) assesses the proportion of variance explained by the model.
Evaluation Results
Table 2 summarizes the predictive performance evaluation of baselines and GeoTransformer models in both GDP and ride-share demand prediction tasks.
GDP Prediction.
As shown in Table 2, all three GeoTransformer models surpass every baseline model. Compared to Random Forest and ViT in panel 1, our best model (linear weighting) outperforms by 27% and 18% respectively in $R^{2}$. In Panel 2, the GAT model exhibits robust decoding capabilities for urban latent representations and delivers the best predictive accuracy among baseline models, while our GeoTransformer with linear weighting surpasses it by 14% in $R^{2}$. This demonstrates that our geospatial attention mechanism effectively models spatial dependencies without the need for predefined graph structures. In panel 3, among the three GeoTransformer models operating under identical parameters and training epochs, the linear weighting approach exhibits the best predictive performance.
Ride-share Demand Prediction.
The results in Panel 1 of Table 2 indicate that ride-share demand is more strongly influenced by sociodemographic factors. Both polynomial regression and GAT models struggle to effectively decode latent representations for accurate predictions. However, our GeoTransformer model with IDW weighting outperforms the best baseline by 18% in $R^{2}$ on the test set. Additionally, among the GeoTransformer variants, it is observed that the IDW model not only achieves the highest performance but also exhibits the fastest convergence.
Figure 4: An example of the attention map between the query region and its neighboring value regions.
Qualitative Analysis.
To investigate the functionality of the geospatial attention mechanism, we visualize several query regions’ attention maps with their neighboring value regions. We present an example in Figure 4, which (a) includes satellite images of the northwest area of the Gresham community, with the query region marked by a red marker (Figure 4), and its attention map for GDP prediction. The query region focuses on its own community area, nearby amenities (the park in the lower left and the supermarket and church in the upper right), and critical road infrastructure (the freeway entrance in the upper left). These results validate the effectiveness of the geospatial attention mechanism in accurately capturing urban features relevant to GDP.
Ablation Study
To investigate the necessity and effectiveness of each design in GeoTransformer, we conduct ablation experiments for geospatial weighting, trainable key matrices, and the number of neighboring value regions.
Figure 5: The GDP prediction accuracy comparison on the test set between no weighting attention and geospatial weighted attention during the training process.
Geospatial Attention Module.
To demonstrate the effectiveness of the geospatial attention module, we conducted a controlled experiment by training an alternative version of the model devoid of the weighting attention mechanism. This baseline model maintained the same architecture and training parameters as the GeoTransformer models, featuring 16 heads, 81 neighbors, 4 layers, and an L2 regularization decay weight of 0.0001, trained over 50 epochs. As shown in Figure 5, the comparative analyses illustrate that all three geospatial attention variants significantly enhanced the GDP prediction accuracy.
Trainable Key Matrices.
Intuitively, the correlation between different regions in urban forecasting is not directly determined by the similarity between these regions. Therefore, we set the key matrix in the attention mechanism as trainable parameters to flexibly adjust the correlation between the value regions and the query region. To validate the effectiveness of this design, we train a GeoTransformer model by setting key matrices $K$ fixed and identical to value matrices $V$, while keeping all other settings constant. As shown in Table 3, linear weighting GeoTransformer with trainable $K$ achieves higher prediction accuracy over 5% in $R^{2}$. It is also observed that the fixed $K$ GeoTransformer converges faster during training, which is reasonable given its smaller number of trainable parameters. However, the flexibility of a trainable $K$ allows the model to ultimately achieve higher accuracy.
[…]
Fixed $K$ 0.241/0.410 0.364/0.464 0.936/0.893
Trainable $K$ 0.081/0.216 0.210/0.346 0.979/0.944
Table 3: Comparison with linear weighting GeoTransformer with fixed $K$ matrix or trainable $K$ matrix.
$k$-Nearest Neighbors.
The spatial extent of surrounding information required by the query region may vary for different urban forecasting tasks. Therefore, we search for the optimal number of neighboring regions $k$ for GDP and ride-share demand predictions. By centering on the query region and expanding outward in a square, $k$ is determined by the square of the side length. We linearly increase the side length to identify the optimal $k$ that maximizes model performance.
For GDP prediction, providing the model with a broader range of urban information initially leads to an improvement in accuracy. However, as $k$ increases further, the model may become overwhelmed by excessive information, resulting in a decline in accuracy (Figure 6).
A similar trend is observed in the ride-share demand prediction, where increasing $k$ initially enhances model accuracy. The model performs best when $k$ is set to 49 or 81. However, further increasing $k$ by incorporating additional neighboring information leads to a decline in accuracy (Figure 6).
Figure 6: Prediction accuracy comparison across $k$ nearest neighbors.
Conclusion and Discussion
We propose a GeoTransformer structure to extract information from high-dimensional urban latent representations. The novel geospatial attention mechanism allows GeoTransformer to capture extensive urban information and its spatial dependency. The experimental results demonstrate the capacity of our method in urban forecasting tasks. An ablation study is conducted to validate the effectiveness of the geospatial attention mechanism and the design of GeoTransformer.
Self-attention module in Transformer (Vaswani 2017) is primarily intended for sequence prediction tasks. In urban forecasting, spatiotemporal prediction tasks also resemble sequence prediction. Although this study does not activate the self-attention module, we retain it in the GeoTransformer as it holds the potential for adapting to spatiotemporal sequence prediction tasks. This merits further exploration in subsequent research.

### v2

@@
-GeoTransformer: Enhancing Urban Forecasting with Geospatial Attention Mechanisms
+GeoTransformer: Enhancing Urban Forecasting with Dependency Retrieval and Geospatial Attention
 Abstract
-Recent advancements have focused on encoding urban spatial information into high-dimensional spaces, with notable efforts dedicated to integrating sociodemographic data and satellite imagery.
-These efforts have established foundational models in this field.
-However, the effective utilization of these spatial representations for urban forecasting applications remains under-explored.
-To address this gap, we introduce GeoTransformer, a novel structure that synergizes the Transformer architecture with geospatial statistics prior.
-GeoTransformer employs an innovative geospatial attention mechanism to incorporate extensive urban information and spatial dependencies into a unified predictive model.
-Specifically, we compute geospatial weighted attention scores between the target region and surrounding regions and leverage the integrated urban information for predictions.
-Extensive experiments on GDP and ride-share demand prediction tasks demonstrate that GeoTransformer significantly outperforms existing baseline models, showcasing its potential to enhance urban forecasting tasks.
+Recent advances in urban forecasting have leveraged high-dimensional spatial data through two primary approaches: graph-based methods that rely on predefined spatial structures and region-based methods that use satellite imagery for local features.
+Although these methods have laid an important foundation, they struggle to integrate holistic urban information and dynamically model spatial dependencies.
+To address this gap, we propose GeoTransformer, a framework combining high-dimensional regional embeddings with dynamic spatial modeling.
+GeoTransformer features two innovations: (1) a dependency retrieval module identifying spatial dependencies to select relevant regions, and (2) a geospatial attention mechanism leveraging global urban information.
+These components unify structural and global urban information for better predictions.
+Extensive experiments on GDP and ride-share demand forecasting show that GeoTransformer outperforms baselines, highlighting its effectiveness in advancing urban forecasting tasks.
 Introduction
-In urban forecasting tasks involving predicting economic indicators and human mobility, classical methods usually rely on low-dimensional numeric data, which includes Point of Interest (POI) data(Li et al.
+In urban forecasting tasks involving predicting economic indicators and human mobility, classical methods usually rely on low-dimensional numeric data, which includes spatial features, socioeconomic data, human mobility data, and so forth (Li et al.
+2022; Roskam and Kunst 2008; DHS et al.
+2013; Wang et al.
+2023; Moreira-Matias et al.
+2013; Pappalardo et al.
+2013).
+Recently, advances in natural language processing (NLP) and deep learning have enabled the use of high-dimensional information to capture the complexities of urban dynamics.
+Some studies have transformed urban information into high-dimensional representations, either through text embedding (Chen et al.
+2022b; Li et al.
+2020, 2023), spatial representation learning (Mai, Li, and Lao 2023; Wu et al.
+2024b) or by encoding satellite imagery data (Jean et al.
+2018; Wang et al.
+2024), offering a new urban modeling paradigm by leveraging high-dimensional data to capture complex urban dynamics.
+High-dimensional urban forecasting applications can be broadly categorized into two directions.
+The first utilizes graph-based modeling with spatial feature embeddings, then using Graph Neural Networks (GNNs) or Graph Attention Networks (GATs) for predictions (Feng and Tassiulas 2022; Zheng et al.
+2020; Chen et al.
+2022b; Li et al.
+2020, 2023).
+While effective, these methods depend heavily on predefined spatial structures and high-quality spatial data, which limits their flexibility in data-sparse or dynamically changing environments.
+Another direction focuses on deriving high-dimensional representations directly from satellite imagery (Jean et al.
+2018; Wang et al.
+2024).
+These methods avoid reliance on predefined structures by encoding urban regional information into independent high-dimensional patches.
+However, these approaches only utilize local information within each patch for prediction and lack the capability to incorporate global urban context (Wang et al.
+2024), which is crucial for tasks requiring holistic understanding.
+To address these limitations, we propose GeoTransformer, a novel transformer-based (Vaswani et al.
+2017) architecture that leverages high-dimensional regional embeddings without requiring predefined spatial structures.
+GeoTransformer incorporates two key innovations: (1) a dependency retrieval module that infers and integrates urban structural dependencies to retrieve the most relevant regions; and (2) a geospatial attention mechanism, utilizing cross-attention to dynamically capture global urban information, enhanced with spatial proximity and information entropy weighting.
+These innovations enable GeoTransformer to effectively combine local and global perspectives, offering a robust framework for diverse urban prediction tasks.
+Extensive experiments demonstrate that our model significantly outperforms baseline methods in predicting GDP and ride-share demand.
+Our main contributions can be summarized as follows:
+  • We propose GeoTransformer, a novel deep architecture that dynamically models high-dimensional urban representations by integrating global spatial dependencies while eliminating the reliance on predefined spatial structures.
+  • We introduce two novel modules: a dependency retrieval module that captures urban structures, and a geospatial attention mechanism that prioritizes representations based on spatial proximity and information entropy.
+  • Experiments demonstrate that our method significantly outperforms baseline models in tasks of GDP and ride-share demand prediction.
+Related Work
+High-Dimensional Representations for Urban Forecasting
+Urban forecasting has increasingly leveraged high-dimensional data to address complex challenges.
+Traditional approaches often rely on low-dimensional numerical data, such as Point of Interest (POI) data(Li et al.
 2022), survey data(Roskam and Kunst 2008; DHS et al.
 2013), GPS records (Wang et al.
-2023a; Moreira-Matias et al.
+2023; Moreira-Matias et al.
 2013; Pappalardo et al.
-2013), demographic census, geospatial features and so forth.
-Recently, many researchers have explored using high-dimensional information to fully capture the complexities of urban dynamics.
-A main direction is utilizing urban imagery for feature extraction and prediction (Ayush et al.
-2020; Li et al.
-2022; Yeh et al.
-2020; Abdelhalim and Zhao 2024).
-A more advanced direction is encoding the urban information to a high-dimensional space representation, either through spatial representation learning (Mai, Li, and Lao 2023; Wu et al.
-2024) or fusing satellite images and sociodemographic data (Wang et al.
-2024).
-While these studies have established frameworks for encoding urban information, there has been limited investigation into optimizing these high-dimensional data for urban forecasting.
-Numerous established methods in natural language processing (NLP) and computer vision have effectively harnessed high-dimensional data.
-Particularly, Transformer architecture and attention mechanisms have proven highly effective in decoding high-dimensional representations of natural language (Vaswani et al.
-2017a; Devlin et al.
-2018; Radford et al.
-2018) and images (Dosovitskiy et al.
-2020a; Rombach et al.
-2022).
-Urban representations, characterized by high dimensionality and spatial correlations, are well-suited for applying similar decoding methods.
-Some studies have explored the use of attention mechanisms to capture spatial dependencies, particularly through Graph Attention Networks (GAT) (Feng and Tassiulas 2022; Zheng et al.
-2020).
-However, these approaches primarily depend on graph structures and spatial embeddings.
-In fact, incorporating geospatial statistical priors might be more effective in encoding spatial dependencies, potentially reducing the need for complex data structuring and simplifying the system design.
-To address this issue, we propose GeoTransformer, a deep architecture akin to Transformer (Vaswani et al.
-2017b) that combines geospatial characteristics prior to the attention mechanism.
-The novel geospatial attention mechanism calculates attention between different urban regions weighted by distance, incorporating information from surrounding areas.
-This approach implicitly includes spatial dependencies in the attention mechanism.
-Extensive experiments demonstrate that our model significantly outperforms several baseline methods in predicting GDP and ride-share demand.
-Our main contributions can be summarized as follows:
-  • We propose GeoTransformer, a transformer-akin deep architecture designed for decoding high-dimensional urban representations and enhancing prediction accuracy.
-  • We introduce a geospatial attention mechanism that prioritizes attention between regions weighted by spatial proximity, assigning higher weights to closer regions.
-  • Experiments demonstrate that our method significantly outperforms baseline models in tasks of GDP and ride-share demand prediction.
-Related Work
-High Dimensional Input for Urban Forecasting
-Urban forecasting leverages multiple data sources and advanced computational techniques to address challenges faced by cities.
-Traditional methods usually adopt low-dimensional numeric data for prediction.
-In recent years, with the need rising for complex prediction tasks and the development of deep learning techniques, researchers have explored various ways to leverage high-dimensional urban information.
-One direction is directly using various urban imagery for feature extraction and prediction.
-Albert, Kaur, and Gonzalez employ Convolutional Neural Networks (CNNs) to detect land-use patterns from satellite images.
-Ayush et al.; Li et al.; Yeh et al.
-utilize urban imagery to predict socioeconomic status.
-Abdelhalim and Zhao leverage roadside imagery to predict transit travel time.
-Recently, encoding urban information into a high-dimensional representation marks a new direction in enhancing flexibility and prediction accuracy for different urban forecasting tasks.
-Mai, Li, and Lao develop spatial representation learning (SRL), utilizing deep neural networks (DNNs) to encode and feature spatial data in the forms of points, polylines, polygons, graphs, etc.
-Wang et al.
-propose a theoretical framework deep hybrid model (DHM) fusing demographic numeric data and satellite images into urban latent representations.
-However, how to leverage these urban representations more effectively in prediction tasks remains unexplored.
-Transformer and attention mechanism(Vaswani 2017) has been widely adopted for its efficiency in capturing relationships within high-dimensional representations.
-They are first proposed to solve natural language understanding and then visual models like ViT(Dosovitskiy et al.
-2020b) and Stable Diffusion (Rombach et al.
-2022) also utilize attention mechanisms to comprehend images and the spatial relationships between image patches.
-Given its strength in handling high-dimensional data and capturing spatial relationships, the attention mechanism is well-suited for adaptation in modeling urban high-dimensional representations.
-Approaches to Modeling Spatial Dependency
-Spatial dependency has traditionally been addressed using two primary approaches.
-First, feature engineering methods like k-nearest neighbors (KNN) are employed to convert surrounding spatial information into features.
-Second, spatial statistical methods such as spatial error regression, spatial lag models(Anselin 2013), and geographically weighted regression (GWR) (Fotheringham, Charlton, and Brunsdon 1998) have been widely used to account for spatial autocorrelation.
-In recent years, Graph Neural Networks (GNNs) have become a widely adopted method for capturing spatial dependencies, especially in spatiotemporal prediction(Wang et al.
-2023b; Zhuang et al.
-2023; Cai et al.
-2020).
-Urban built environment data such as road networks, are inherently suitable for representation as graph structures due to their topological properties.
-However, grid-based high-dimensional urban representations often lack clear graph-like structures, making graph construction particularly challenging.
-Moreover, the intrinsic complexity of such high-dimensional data can further hinder the effective training of models.
-Recent advancements have seen the integration of graph structures with attention mechanisms (Zheng et al.
-2020; Feng and Tassiulas 2022), particularly through Graph Attention Networks (GATs)(Zhang, Yu, and Liu 2019).
-While these models incorporate spatiotemporal attention mechanisms, their reliance on graph structures and spatial embeddings remains significant for capturing spatial dependencies.
-Integrating spatial statistical priors could simplify the model’s reliance on complex data structures and streamline system design.
-Tobler’s First Law of Geography (Tobler 1970), which posits that closer entities are more likely to be related, could provide a valuable framework for capturing spatial dependency.
-Despite its potential, the integration of spatial statistical priors with attention mechanisms remains unexplored.
-Therefore, in this study, we aim to introduce an attention mechanism guided by spatial statistical priors as a new paradigm for effectively leveraging high-dimensional urban representation data.
+2013), demographic census, spatial features and so forth, limiting their ability to capture the multifaceted nature of urban dynamics.
+Recent advances have focused on two primary directions for high-dimensional urban representations.
+The first direction employs graph-based methods that embed urban information into representations by constructing predefined graph structures.
+Graph Attention Networks (GATs) (Zhang, Yu, and Liu 2019) are used to infer POI relationships (Chen et al.
+2022b; Li et al.
+2020, 2023).
+Feng and Tassiulas propose an Adaptive Graph Spatial-Temporal Transformer Network to model cross-spatial-temporal correlations.
+Zheng et al.
+use a graph multi-attention network to model the impact of the spatio-temporal factors on traffic conditions.
+Chen et al.
+propose a spatial-aware attention module based on spatial proximity.
+Additionally, position embeddings are employed as learnable parameters to provide location information (Xu et al.
+2021).
+These methods rely heavily on predefined spatial structures, such as road networks or adjacency matrices, which limits their applicability in dynamic or data-sparse urban environments.
+The second direction focuses on region-based methods that leverage satellite imagery and other high-resolution data to encode urban areas into high-dimensional representations.
+Tile2Vec (Jean et al.
+2018) learns representations from satellite imagery tiles.
+(Wang et al.
+2024) propose a deep hybrid model fusing regional built environment and socio-demographic information into latent representations through multi-task learning.
+These region-based representation learning methods avoid reliance on predefined graph structures.
+However, they only leverage local information patches for prediction (Wang et al.
+2024) and struggle to incorporate global urban context or dynamically model spatial dependencies.
+Additionally, region-based representations face variability in effectiveness due to encoding uncertainty, which remains unaddressed.
+However, information entropy has been shown to evaluate the effectiveness of high-dimensional data (Wu et al.
+2024a) and has been used for weighting (Zhu, Tian, and Yan 2020).
+It offers a method to mitigate these disparities among region-based urban representations, presenting a promising yet underexplored solution.
+Urban Spatial Dependency Modeling
+Spatial dependency provides key urban structural insights.
+It has traditionally been addressed through two main approaches for low-dimensional numerical data: feature engineering methods, such as k-nearest neighbors (KNN), and spatial statistical techniques, including spatial error regression, spatial lag models (Anselin 2013), and geographically weighted regression (GWR) (Fotheringham, Charlton, and Brunsdon 1998).
+For high-dimensional representations, GATs employ attention mechanisms to aggregate information from neighboring nodes (Feng and Tassiulas 2022; Zheng et al.
+2020; Chen et al.
+2022b; Xu et al.
+2021).
+However, region-based high-dimensional urban representations often lack explicit graph-like structures, further complicating their ability to capture spatial dependencies.
+Data-driven methods like sparse linear layers may provide insights to dynamically capture spatial dependency (Li and Moura 2020), yet their applications on region-based spatial representations remain unexplored.
+Therefore, to address these challenges, we propose leveraging a sparse matrix-based dependency retrieval mechanism to dynamically model urban structures and integrating geospatial cross-attention weighted by spatial proximity and information entropy to incorporate both local and global information.
+These techniques aim to overcome the limitations of predefined graph structures and local-only region-based approaches in existing studies.
 Economic and Mobility Indicators Prediction
@@
 There is substantial research indicating a strong association between GDP and urban spatial characteristics (Zhang, He, and Yuan 2020; Zhaoa and Chenb 2023), with numerous studies employing machine learning methods to model and predict regional GDP using structured urban information or urban imagery data(Kopczewska 2022; Chen et al.
-2022b).
+2022c).
 For example, Chen et al.
@@
 2023; Zhao et al.
-2022; Zhang and Zhao 2022).
+2022; Tu et al.
+2024).
 Zhao et al.
@@
 Methodology
-As illustrated in Figure 1, our framework adopts an encoder-decoder structure.
-The encoder leverages existing state-of-the-art methods to process numeric and imagery inputs, representing city regions in a latent space through a mixing operator.
-Our novel GeoTransformer decoder model then utilizes this latent representation to predict GDP and ride-share demand.
+GeoTransformer
+As illustrated in Figure 1, the whole framework adopts an encoder-decoder structure.
+The encoder leverages a region-based representation method to process numeric and imagery inputs, representing city regions in a latent space through a mixing operator.
+We propose a novel approach GeoTransformer to decode the representations for downstream tasks including GDP and ride-share demand prediction.
+Our method contains two key innovations: a dependency retrieval module and a geospatial attention module.
 Figure 1: Our urban forecasting framework uses a mixing operator to fuse information into latent representations which are then passed to GeoTransformer for prediction.
-Encode the city: Mixing Operator
-To fuse satellite imagery and sociodemographic information into a high-dimensional latent representation, we adopt a mixing operator $\mathcal{M}$, which can be represented as:
-$z_{n}=\mathcal{M}(x_{n},I_{n})$ (1)
-in which $x_{n}$ and $I_{n}$ represent the sociodemographics and satellite imagery inputs, and $z_{n}$ represents the latent representation of an urban area.
-Following the work Deep Hybrid Model (Wang et al.
-2024), we adopt a supervised autoencoder (SAE).
-The SAE maps the imagery into latent variables through the encoder $E(I_{n})$, and reconstructs images through the decoder $D(z_{n})$ and sociodemographic through a supervision network $F(z_{n})$ (Figure 2 ).
-It is trained by optimizing the sum of autoencoder loss $\mathcal{L}_{AE}$ and sociodemographic supervision loss $\mathcal{L}_{sup}$:
-$\mathcal{L}_{SAE}=(1-\lambda)\mathcal{L}_{AE}+\lambda\mathcal{L}_{sup}$ (2)
-where $\mathcal{L}_{sup}$ denotes the absolute error of the sociodemographic variables $x$ $\mathcal{L}_{AE}$ represents the sum of L1 image reconstruction loss $\mathcal{L}_{rec}$, the KL divergence loss $\mathcal{L}_{kl}$, and the learned perceptual image patch similarity loss $\mathcal{L}_{lpipis}$:
-$\mathcal{L}_{sup}=\alpha_{0}|x-\hat{x}|$ (3)
-$\mathcal{L}_{AE}=\alpha_{1}\mathcal{L}_{rec}+\alpha_{2}\mathcal{L}_{kl}+\alpha_{3}\mathcal{L}_{lpipis}$ (4)
-in which all the $\alpha$ terms are the adaptive hyperparameters used for scaling each loss.
-For each autoencoder loss, $\mathcal{L}_{rec}$ measures the L1 distance between the input images $I_{n}$ and reconstructed images $\hat{I_{n}}$ in the pixel space, $\mathcal{L}_{kl}$ measures the distance between the estimated latent distribution and the standard normal prior distribution, and $\mathcal{L}_{lpips}$ is calculated by extracting feature stacks from $L$ layers of a pre-trained neural network, unit-normalizing them in the channel dimension ($\hat{y}^{l}_{I_{n,hw}},\hat{y}^{l}_{\hat{I}_{n,hw}}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$), and scaling by vector $w_{l}$ before computing the L2 distance.
-The mixing operator allows for the fusion of sociodemographics and satellite imagery information in a high-dimensional latent space.
+As shown in Figure 2, our method adopts a structure akin to the Transformer (Vaswani et al.
+2017) decoder.
+Most relevant regions are retrieved from all urban regions.
+Then the target and retrieved regions are encoded into representations and then input into the transformer decoder for prediction.
+The GeoTransformer system $\mathcal{G}$ can be represented as:
+$y_{i}=\mathcal{G}(z_{i},\text{Retrieval}(i,Z))$ (1)
+in which the input includes the latent representation $z_{i}$ of the target region $i$ and retreived representations $Z_{i}=\text{Retrieval}(i,Z)$ from all regions $Z$, and $y_{i}$ is the predicted output.
+Figure 2: The GeoTransformer framework includes a retrieval module and a geospatial attention module that calculates attention scores between the query region and retrieved value regions.
+Region-based Urban Representations
+To obtain the regional representations, we adopt a mixing operator $\mathcal{M}$ following the work Deep Hybrid Model (Wang et al.
+2024), to fuse satellite imagery and sociodemographic information into a high-dimensional latent space, which can be represented as:
+$z_{i}=\mathcal{M}(x_{i},I_{i})$ (2)
+in which $x_{i}$ and $I_{i}$ represent the sociodemographics and satellite imagery inputs, and $z_{i}$ represents the latent representation of an urban area.
+The supervised autoencoder (SAE) is deployed in this framework to learn the representations through optimizing an autoencoder loss and a sociodemographic supervision loss (Figure 3).
 Each region of a city is then encoded into a latent representation using the encoder of the trained SAE.
 Figure 3: The mixing operator is trained by extracting sociodemographic information and reconstructing the input satellite images from latent space.
-GeoTransformer: Geospatial-weighted Attention Mechanisms
-We propose a novel approach GeoTransformer to decode the urban latent representations for downstream tasks including GDP and ride-share demand prediction.
-As shown in Figure 3, our method adopts a structure akin to the Transformer (Vaswani 2017) decoder.
-Since the downstream tasks do not involve sequence prediction, the self-attention module is optional in this study.
-We introduce a geospatial attention mechanism to capture the spatial dependencies and provide location information between city regions.
-The outputs of the decoder layers are then passed through a fully connected linear layer.
-The whole system $\mathcal{G}$ can be represented as:
-$y_{n}=\mathcal{G}(z_{n},\text{kNN}(z_{n},Z))$ (5)
-in which the input includes the latent representation of a region $z_{n}$ and its k-nearest neighbors’ latent representations $\text{kNN}(z_{n},Z)$ from all regions $Z$, and $y_{n}$ is the predicted output.
-The geospatial attention mechanism is the core component of our method.
-It computes attention scores between city regions and applies geospatial weighting to these scores.
+Dependency Retrieval Module
+To address the challenge of modeling dynamic inter-region interactions for region-based representation methods, we employ a sparse matrix to explicitly represent spatial dependencies.
+The sparse matrix encodes the influence of one region on another through weights, where only a subset of regions significantly influence the given target region.
+Building on this, we propose a dependency retrieval mechanism that dynamically identifies the most relevant regions around the target region as the GeoTransformer model inputs.
+Specifically, for each target region, the mechanism first identifies the $k_{1}$-nearest neighbors based on spatial proximity.
+Then, from these $k_{1}$ candidates, it selects the top-$k_{2}$ regions with the highest weights in the sparse matrix.
+The retrieval mechanism is defined as:
+$Z_{i}=\{z_{j}\mid j\in\mathrm{arg\,top}_{k_{2}}(S[i,\mathcal{N}_{k_{1}}(i)])\},$ (3)
+where $Z_{i}$ denotes the set of high-dimensional representations $z_{j}\in\mathbb{R}^{d}$ for the $k_{2}$ most relevant regions selected from the $k_{1}$-nearest neighbors of the target region $i$.
+Here, $\mathcal{N}_{k_{1}}(i)$ represents the indices of the $k_{1}$-nearest neighbors of region $i$, identified based on spatial proximity.
+$S[i,j]$ is the weight from the sparse matrix $S$, which quantifies the dependency of region $i$ on region $j$.
+The sparse matrix $S$ is constructed using Lasso regression, which enforces sparsity to identify the most important inter-region dependencies.
+For a given target region $i$, its high-dimensional representation $z_{i}\in\mathbb{R}^{d}$ is treated as the target variable, while the representations of all other regions $X$ serve as input features.
+The optimization objective for Lasso is defined as:
+$\min_{\beta}\|z_{i}-X\beta\|_{2}^{2}+\alpha\|\beta\|_{1},$ (4)
+where $X\in\mathbb{R}^{(n-1)\times d}$ denotes the matrix of feature representations for all other regions except $i$.
+The sparse weight vector $\beta\in\mathbb{R}^{n-1}$ captures the influence of other regions on region $i$, while $\alpha$ is the regularization strength that controls the degree of sparsity in the resulting weights.
+The Lasso regularization term $\|\beta\|_{1}$ forces most weights to zero, retaining only the strongest dependencies.
+The resulting sparse weight vector $\beta$ is then used to populate the corresponding row $S[i,:]$ in the sparse matrix $S$.
+This dependency retrieval module enables region-based representations to effectively model spatial dependency without predefined urban structure information, providing a foundation for attention computing in the GeoTransformer model.
+Geospatial Attention Module
+We introduce a geospatial attention mechanism to leverage the global urban information for prediction.
+As shown in Figure 2, the model first calculates cross-attention between the target and retrieved representations, with the attention scores weighted by spatial proximity and information entropy.
 The attention calculation is represented as:
-$\text{GeoAtt}(Q,K,V)=W\odot\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (6)
-where the query matrix $Q$ contains the latent representation of the query region $z_{n}$ for which the target variable is being predicted, while the value matrix $V$ consists of $z_{n}$ itself and the latent representations of its neighbors as value regions.
+$\text{GeoAtt}(Q,K,V)=(\lambda W_{s}+(1-\lambda)W_{e})\odot\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (5)
+where the query matrix $Q$ contains the representation of the query region $z_{i}$ for which the target variable is being predicted, while the value matrix $V$ consists of $z_{i}$ itself and the retrieved representations as value regions.
 $d_{k}$ denotes the dimension of the key vectors.
-Intuitively, as the similarity between two city regions is not directly related to target variables, we set the key matrix $K$ as trainable weights rather than being identical to $V$.
-$W$ is the spatial weighting factor, computed based on the spatial distance between query regions in $Q$ and value regions in $V$.
-Our model provides linear and non-linear spatial weighting options based on the assumption of Tobler’s First Law of Geography (Tobler 1970), suggesting that regions closer to the query region are assigned higher weights.
-Linear Normalization
-scales the distances within 0 to 1, inversely transforming shorter distances into higher weights:
-$W_{j}=1-\frac{d_{j}}{\text{max}(d)}$ (7)
+Intuitively, as the similarity between two city regions is not directly related to prediction variables, we set the key matrix $K$ as trainable weights rather than being identical to $V$.
+$W_{s}$ and $W_{e}$ are the spatial proximity and information entropy weighting factors respectively.
+Their relative importance is controlled by a hyperparameter $\lambda\in[0,1]$.
+Spatial Proximity Weighting
+Our model provides spatial proximity weighting as a way to provide location information.
+Based on the assumption of Tobler’s First Law of Geography (Tobler 1970), we assigned regions closer to the query region with higher weights.
+We leverage a linear weighting method, scaling the distances within 0 to 1, inversely transforming shorter distances into higher weights:
+$W_{S_{j}}=1-\frac{d_{j}}{\text{max}(d)}$ (6)
 where $d_{j}$ is the distance to the $j^{th}$ value region, and $\text{max}(d)$ are the maximum distances observed.
-Inverse Distance Weighting (IDW) (Shepard 1968)
-calculates weights inversely proportional to the distance:
-$W_{j}=\frac{1}{d^{p}_{j}+\epsilon}$ (8)
-where $d_{j}$ is the distance to the $j^{th}$ value region, the power value $p$ is set to 2 in standard applications, and $\epsilon$ is a small constant to prevent division by zero.
-The weights are then subsequently normalized to ensure they sum to unity.
-Gaussian Weighting
-applies a Gaussian function to the distances, providing a gradual decrease in weight as distance increases:
-$W_{j}=e^{-(\frac{d^{2}_{j}}{2\sigma^{2}})}$ (9)
-where $d_{j}$ is the distance to the $j^{th}$ value region, and $\sigma$ determines the spread of the weights. The weights are then subsequently normalized to ensure they sum to unity.
-Adhering to the established multi-head attention mechanism (Vaswani 2017), our model employs multiple attention heads to capture various spatial relationships concurrently.
+Information Entropy Weighting
+As a way to evaluate the effectiveness of high-dimensional data, we assign higher weights for representations with higher information entropy.
+Regions with higher entropy are believed to contain richer and more complex urban information, thereby playing a more significant role in prediction.
+The calculation can be presented as:
+$W_{E_{j}}=\frac{H_{j}}{\max(H_{1},H_{2},\dots,H_{n})}$ (7)
+where $H_{i}$ is the Shannon entropy of the latent representation $z_{i}$, and $\max(H_{1},H_{2},\dots,H_{n})$ is the maximum entropy across retrieved tiles.
+Since the the latent representation is a high-dimensional vector $z_{i}$, the entropy $H_{i}$ for each individual latent representation cab be represented as:
+${H}_{j}=-\sum_{i=1}^{d}p_{ji}\log(p_{ji})$ (8)
+where $\textbf{where }\quad\text{p}_{ij}=\frac{e^{z_{ij}}}{\sum_{k=1}^{d}e^{z_{ik}}}$ (9)
+Here, $z_{ji}$ is the value of the $j-{th}$ feature for latent representation $z_{i}$, and $p_{ij}$ is the normalized probability.
+Adhering to the established multi-head attention mechanism (Vaswani et al.
+2017), our model employs multiple attention heads to capture various spatial relationships concurrently.
 The calculation can be represented as:
@@
 $\textbf{where }\text{head}_{i}=\text{GeoAtt}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i})$ (11)
-where $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$ and $W^{O}$ are the projection matrices. The amount of heads and layers of the decoder module is also adaptive.
-The GeoTransformer architecture is designed to extract geospatial information in high-dimensional latent space, and it is also adaptable to various geospatial prediction tasks due to the flexibility of the output layers.
-The design of geospatial attention allows the model to capture and integrate spatial dependencies, utilizing local context from both the target region and its surroundings for more accurate predictions.
-Leveraging high-dimensional urban information and deep architectures, our GeoTransformer represents a novel approach to geospatial analysis and urban planning applications.
-Figure 3: The GeoTransformer structure includes a geospatial attention module that calculates the geospatial weighted attention scores between the query region and its neighboring value regions.
-Panel 1 Baselines with only numeric or imagery input
-Model 1: Linear Regression sociodemographic numeric input (12 dim)
-Model 2: ViT satellite imagery input (4*512*512 dim), heads = 12, layers = 12
-Panel 2 Baselines with urban latent representations input
-Model 3: Polynomial Regression latent input (reduced dim from 4096 to 128 using PCA), poly degree = 2, $l_{2}$ regularization with $\lambda$ = 0.5
-Model 4: GAT latent input (4096 dim), layers = 4
-Panel 3 GeoTransformer with 3 geospatial weighting methods
-Model 5: Linear Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=49 / 81
-Model 6: IDW Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=49 / 81
-Model 7: Gaussian Weighting latent input (4096 dim), heads = 16, layers = 4, neighbors=81 / 169
+where $W^{Q}_{i}$, $W^{K}_{i}$, $W^{V}_{i}$ and $W^{O}$ are the projection matrices.
+The amount of heads and layers of the decoder module is also adaptive.
+At last, the outputs of the decoder layers are passed through a fully connected linear layer.
+The GeoTransformer framework is designed to overcome limitations in region-based urban representation methods, capturing spatial dependency dynamically and leveraging global urban information for prediction.
+It is also adaptable to various urban prediction tasks due to the flexibility of the output layers.
+Overall, our GeoTransformer represents a novel approach to geospatial analysis and urban planning applications.
+Panel 1 Deep Hybrid Model Framework (Latent Representations + Regression)
+Model 1: DHM (linear) representation input (4096 dim), Linear Regression
+Model 2: DHM (polynomial) representation input (reduced dim from 4096 to 128 using PCA), Polynomial Regression (poly degree = 2, $l_{2}$ regularization $\lambda$ = 0.5)
+Panel 2 Latent Representations + Graph Attention Networksl
+Model 3: GAT (grid) representation input (4096 dim), layers = 4
+Model 4: GAT (sparse) representation input (4096 dim), layers = 4, retrieval size = 81
+Panel 3 Latent Representations + GeoTransformer (Ours)
+Model 5: GeoTransformer representation input (4096 dim), heads = 16, layers = 4, KNN=121, retrieval size = 81
 Table 1: Training settings of baselines and GeoTransformer models.
@@
 MSE $\downarrow$ MAE $\downarrow$ $R^{2}$ MSE×$\times 10^{7}$downarrow MAE×$\times 10^{3}$downarrow $R^{2}$
-Panel 1 Baselines with only numeric or imagery input
-1 Linear Regression 3.22/3.13 1.44/1.43 0.159/0.151 8.45/6.11 8.36/5.99 0.668/0.720
-2 ViT 0.616/0.828 0.595/0.703 0.822/0.759 15.6/22.5 4.10/4.16 0.526/0.315
-Panel 2 Baselines with urban latent representations input
-3 Polynomial Regression 2.19/3.25 1.17/1.46 0.429/0.118 12.8/19.5 5.05/6.00 0.494/0.335
-4 GAT 0.469/0.726 0.476/0.602 0.866/0.805 7.21/9.51 2.01/2.35 0.750/0.575
-Panel 3 GeoTransformer with 3 geospatial weighting methods
-5 Linear Weighting 0.081/0.216 0.210/0.346 0.979/0.944 1.62/7.26 0.98/2.53 0.934/0.779
-6 IDW Weighting 0.570/0.686 0.537/0.585 0.849/0.822 0.172/3.02 0.491/1.16 0.993/0.908
-7 Gaussian Weighting 0.448/0.567 0.486/0.533 0.882/0.852 0.263/9.78 0.469/3.97 0.989/0.703
+Panel 1 Deep Hybrid Model Framework (Latent Representations + Regression)
+1 DHM (linear) 3.56/3.23 1.84/1.75 0.151/0.142 9.03/7.28 9.36/8.90 0.568/0.601
+2 DHM (polynomial) 2.19/3.25 1.17/1.46 0.429/0.118 12.8/19.5 5.05/6.00 0.494/0.335
+Panel 2 Latent Representations + Graph Attention Networks
+3 GAT (grid) 0.469/0.726 0.476/0.602 0.866/0.805 7.21/9.51 2.01/2.35 0.750/0.575
+4 GAT (sparse) 0.592/0.691 0.547/0.561 0.852/0.812 6.22/5.47 1.66/1.83 0.781/0.794
+Panel 3 Latent Representations + GeoTransformer (Ours)
+5 GeoTransformer 0.079/0.216 0.207/0.343 0.964/0.933 0.217/2.94 0.684/1.02 0.920/0.901
 Table 2: Predictive performance of baselines and GeoTransformer models. Each entry is represented as training/testing performance.
 Experiment
-In this section, we provide our method’s results in GDP and ride-share demand prediction and compare them with baseline methods.
-Ablation experiments are conducted to demonstrate the effectiveness of the geospatial attention component and model architecture design.
+In this section, we provide our method’s experiments in GDP and ride-share demand prediction and compare them with baseline methods.
+Ablation experiments are conducted to demonstrate the effectiveness of the dependency retrieval module and two weighting methods in the geospatial attention module.
 Experiment Setup
@@
 We train various baseline models for comparison.
-As detailed in Table 1, models in Panel 1 utilize either direct numeric or imagery input.
-Specifically, Model 1 employs a linear regression configured with 12-dimensional sociodemographic data as input.
-Model 2 utilizes a pre-trained Visual Transformer (ViT) (Dosovitskiy et al.
-2020a), which is subsequently fine-tuned on the prediction tasks with the last 4 encoder layers and the output layer unfrozen.
-Panel 2 presents models that utilize latent representations as inputs to perform predictions.
-Following DHM(Wang et al.
-2024), model 3 employs a polynomial regression.
-Model 4 adopts a Graph Attention Network (GAT) (Zhang, Yu, and Liu 2019) model with each city region as a node connected to its neighbors, forming the graph structure.
-All hyperparameters for training the baseline models are optimized through a grid search to ensure peak performance on the test set.
+As detailed in Table 1, models in Panel 1 utilize the deep hybrid model (DHM) (Wang et al.
+2024) framework, directly inputting representations to regression models to predict.
+Specifically, Model 1 employs a linear regression, and Model 2 utilizes a polynomial regression.
+Panel 2 presents models that utilize representations as inputs for graph attention networks (GATs).
+Model 3 forms the graph structure by assigning each region as a node connected to its rook neighbors.
+Model 4 constructs a graph using the sparse matrix, where each region is a node connected to its top-$k$ relevant neighbors selected by the highest weights in our sparse matrix, with edges weighted accordingly.
+The number of top-$k$ connections in Model 4 is consistent with the retrieval size in GeoTransformer, while all other hyperparameters for training the baseline models are optimized through a grid search to ensure peak performance on the test set.
 GeoTransformer Training.
-As shown in Table 1 panel 3, we train the GeoTransformer (model 5-7) using three spatial weighting methods.
-The number of decoder layers is set to 4, and the number of attention heads is set to 16.
-For each model variant and task, the L2 regularization decay weight was optimized, with a range of 0.0001 to 0.1, while the number of neighboring regions is set to 49, 81, or 169.
+As shown in Table 1 panel 3, we train the GeoTransformer (model 5) with the number of decoder layers set to 4, and the number of attention heads set to 16.
+The L2 regularization decay weight is optimized at 0.001, while the number of retrieved regions is set to 81.
 All parameters are grid-searched to ensure the best performance.
@@
 Evaluation Results
-Table 2 summarizes the predictive performance evaluation of baselines and GeoTransformer models in both GDP and ride-share demand prediction tasks.
+Table 2 summarizes the predictive performance evaluation of baselines and the GeoTransformer model in both GDP and ride-share demand prediction tasks.
 GDP Prediction.
-As shown in Table 2, all three GeoTransformer models surpass every baseline model.
-Compared to Random Forest and ViT in panel 1, our best model (linear weighting) outperforms by 27% and 18% respectively in $R^{2}$.
-In Panel 2, the GAT model exhibits robust decoding capabilities for urban latent representations and delivers the best predictive accuracy among baseline models, while our GeoTransformer with linear weighting surpasses it by 14% in $R^{2}$.
-This demonstrates that our geospatial attention mechanism effectively models spatial dependencies without the need for predefined graph structures.
-In panel 3, among the three GeoTransformer models operating under identical parameters and training epochs, the linear weighting approach exhibits the best predictive performance.
+As shown in Table 2, our GeoTransformer model surpasses every baseline model.
+For Deep Hybrid Models in panel 1, linear regression performs a little better than polynomial regression, and our model outperforms them by around 80% in $R^{2}$.
+In Panel 2, the GAT models exhibit similar capabilities for decoding region-based representations to predict GDP, with over 80% in $R^{2}$, while our GeoTransformer still surpasses them by 10% in $R^{2}$.
+[…]
+1 Random Regions + No Attention Weighting 3.24/3.84 1.54/1.65 0.061/0.007
+2 Retrieved Module + No Attention Weighting 0.618/0.650 0.596/0.708 0.811/0.758
+3 Retrieved Module + Spatial Proximity Weighting 0.245/0.418 0.364/0.478 0.906/0.881
+4 Retrieved Module + Spatial Proximity & Information Entropy Weighting 0.079/0.216 0.207/0.343 0.964/0.933
+Table 3: Ablation experimental results of each GeoTransformer component.
 Ride-share Demand Prediction.
-The results in Panel 1 of Table 2 indicate that ride-share demand is more strongly influenced by sociodemographic factors.
-Both polynomial regression and GAT models struggle to effectively decode latent representations for accurate predictions.
-However, our GeoTransformer model with IDW weighting outperforms the best baseline by 18% in $R^{2}$ on the test set.
-Additionally, among the GeoTransformer variants, it is observed that the IDW model not only achieves the highest performance but also exhibits the fastest convergence.
-Figure 4: An example of the attention map between the query region and its neighboring value regions.
-Qualitative Analysis.
-To investigate the functionality of the geospatial attention mechanism, we visualize several query regions’ attention maps with their neighboring value regions.
-We present an example in Figure 4, which (a) includes satellite images of the northwest area of the Gresham community, with the query region marked by a red marker (Figure 4), and its attention map for GDP prediction.
-The query region focuses on its own community area, nearby amenities (the park in the lower left and the supermarket and church in the upper right), and critical road infrastructure (the freeway entrance in the upper left).
-These results validate the effectiveness of the geospatial attention mechanism in accurately capturing urban features relevant to GDP.
+The results in Panel 1 of Table 2 show that both Deep Hybrid Models achieve 60% and 33% in $R^{2}$ respectively, while GeoTransformer surpasses them over 33 % in $R^{2}$.
+In panel 2, the GAT model based on our sparse matrix achieves nearly 80% in $R^{2}$, around 22% higher than the one with grid structure.
+However, our GeoTransformer model still outperforms it by 13% in $R^{2}$ on the test set.
+Our method significantly outperforms the deep hybrid models relying solely on single-region representations across both prediction tasks, demonstrating the substantial accuracy gains achieved by leveraging global urban information.
+Notably, GeoTransformer’s retrieval mechanism, which captures urban structural information without relying on predefined urban structures, exhibits clear superiority over GAT models that utilize graph structures built from the same information.
+This highlights the effectiveness of GeoTransformer in dynamically modeling urban dependencies and extracting meaningful global context for enhanced predictive performance.
+Qualitative Analysis
+To investigate the functionality of the dependency retrieval module, we visualize several top retrieval results for the target region.
+Figure 4 presents an example of a satellite image of the target region in the northwest area of the Auburn Gresham community and six top relevant regions based on the sparse matrix.
+The relevant regions span two major urban area types.
+The first includes functional zones near the target, such as local middle schools, parks, and churches (Region a-d in Figure 4).
+The second type of area covers transportation hubs within the community, such as railway hub centers and main roads within the community (Region e and f in Figure 4).
+These results showcase our retrieval module’s ability to capture urban structures within the specific spatial scale using the sparse matrix.
+In addition to balancing the inconsistency in the information effectiveness of individual representation samples during the encoding process, we also argue that higher information entropy represents more complex and diverse regions, which is demonstrated by comparing regional representations with high and low information entropy (figure 5).
+Figure 4: An example of dependency retrieval results.
+Figure 5: A comparison of high and low information entropy regions.
 Ablation Study
-To investigate the necessity and effectiveness of each design in GeoTransformer, we conduct ablation experiments for geospatial weighting, trainable key matrices, and the number of neighboring value regions.
-Figure 5: The GDP prediction accuracy comparison on the test set between no weighting attention and geospatial weighted attention during the training process.
-Geospatial Attention Module.
-To demonstrate the effectiveness of the geospatial attention module, we conducted a controlled experiment by training an alternative version of the model devoid of the weighting attention mechanism.
-This baseline model maintained the same architecture and training parameters as the GeoTransformer models, featuring 16 heads, 81 neighbors, 4 layers, and an L2 regularization decay weight of 0.0001, trained over 50 epochs.
-As shown in Figure 5, the comparative analyses illustrate that all three geospatial attention variants significantly enhanced the GDP prediction accuracy.
-Trainable Key Matrices.
-Intuitively, the correlation between different regions in urban forecasting is not directly determined by the similarity between these regions.
-Therefore, we set the key matrix in the attention mechanism as trainable parameters to flexibly adjust the correlation between the value regions and the query region.
-To validate the effectiveness of this design, we train a GeoTransformer model by setting key matrices $K$ fixed and identical to value matrices $V$, while keeping all other settings constant.
-As shown in Table 3, linear weighting GeoTransformer with trainable $K$ achieves higher prediction accuracy over 5% in $R^{2}$.
-It is also observed that the fixed $K$ GeoTransformer converges faster during training, which is reasonable given its smaller number of trainable parameters.
-However, the flexibility of a trainable $K$ allows the model to ultimately achieve higher accuracy.
-[…]
-Fixed $K$ 0.241/0.410 0.364/0.464 0.936/0.893
-Trainable $K$ 0.081/0.216 0.210/0.346 0.979/0.944
-Table 3: Comparison with linear weighting GeoTransformer with fixed $K$ matrix or trainable $K$ matrix.
-$k$-Nearest Neighbors.
-The spatial extent of surrounding information required by the query region may vary for different urban forecasting tasks.
-Therefore, we search for the optimal number of neighboring regions $k$ for GDP and ride-share demand predictions.
-By centering on the query region and expanding outward in a square, $k$ is determined by the square of the side length.
-We linearly increase the side length to identify the optimal $k$ that maximizes model performance.
-For GDP prediction, providing the model with a broader range of urban information initially leads to an improvement in accuracy.
-However, as $k$ increases further, the model may become overwhelmed by excessive information, resulting in a decline in accuracy (Figure 6).
-A similar trend is observed in the ride-share demand prediction, where increasing $k$ initially enhances model accuracy.
-The model performs best when $k$ is set to 49 or 81.
-However, further increasing $k$ by incorporating additional neighboring information leads to a decline in accuracy (Figure 6).
-Figure 6: Prediction accuracy comparison across $k$ nearest neighbors.
+To investigate the necessity and effectiveness of each design in GeoTransformer, we conduct ablation experiments for the dependency retrieval module and two weighting methods for the geospatial attention module.
+The experiment setup and results for GDP prediction are shown in Table 3.
+Four models were designed for the ablation study to evaluate the effectiveness of each component in GeoTransformer.
+Model 1 uses randomly selected regions without any attention weighting, serving as a baseline to demonstrate the importance of retrieval and attention weighting mechanisms.
+Model 2 applies the dependency retrieval mechanism on Model 1 to select relevant regions.
+Model 3 builds upon Model 2 by introducing spatial proximity weighting.
+Finally, Model 4 represents the complete GeoTransformer design, integrating the dependency retrieval module and both spatial proximity weighting and information entropy weighting to refine the attention mechanism.
+The experimental results, presented in Table 3, show a clear and consistent improvement in performance with each added component.
+Model 1, with random selection, yields an $R_{2}$ near zero, indicating a chaotic model, while Model 2 introduces spatial dependencies through the retrieval mechanism, achieving a dramatic leap to over 80% $R_{2}$, demonstrating the effectiveness of the retrieval mechanism in selecting relevant regions.
+The addition of spatial proximity weighting in Model 3 further enhances accuracy, validating the importance of incorporating spatial relationships into the attention module.
+Finally, Model 4 achieves the best performance across all metrics, confirming the necessity of applying information entropy weighting.
 Conclusion and Discussion
-We propose a GeoTransformer structure to extract information from high-dimensional urban latent representations.
-The novel geospatial attention mechanism allows GeoTransformer to capture extensive urban information and its spatial dependency.
-The experimental results demonstrate the capacity of our method in urban forecasting tasks.
-An ablation study is conducted to validate the effectiveness of the geospatial attention mechanism and the design of GeoTransformer.
-Self-attention module in Transformer (Vaswani 2017) is primarily intended for sequence prediction tasks.
-In urban forecasting, spatiotemporal prediction tasks also resemble sequence prediction.
-Although this study does not activate the self-attention module, we retain it in the GeoTransformer as it holds the potential for adapting to spatiotemporal sequence prediction tasks.
-This merits further exploration in subsequent research.
+GeoTransformer demonstrates robust capabilities in urban forecasting, establishing a novel paradigm for region-based representation methods without reliance on predefined urban structural information.
+The experimental results validate the capacity of GeoTransformer, with ablation studies further highlighting the effectiveness of its key components.
+GeoTransformer still has limitations as it focuses solely on static validation and does not explore time-series tasks.
+However, the combination of the novel retrieval mechanism and attention-based model lays a strong foundation for future advancements in spatiotemporal forecasting.